OpenStudy (anonymous):

check my work? 4/3s - 1 = 1/3 a. s= -1 b. s= 0 c. s= 1 d. s= 4/3 i got s = 0 is this correct, if not can you tell me which is?

OpenStudy (uri):

Show 'em work.

OpenStudy (anonymous):

its too hard to type all the fraction work and understand it on here. but i got s=0 is that correct?

OpenStudy (anonymous):

s=1 fits

OpenStudy (anonymous):

just click on the equation button when you are replying and choose fractions from there

OpenStudy (anonymous):

by the way s=1 fits if your equation is like this\[\frac{ 4 }{3s}-1=\frac{ 1}{ 3 }\]

OpenStudy (anonymous):

but if it is \[\frac{ 4 }{3s-1 }=\frac{ 1 }{3 }\] then (3s-1)=12 and s=13/3

OpenStudy (anonymous):

can you show me the work of how you got that?

OpenStudy (anonymous):

sure \[\frac{ 4 }{3s}=1+\frac{ 1 }{3 }\] implies 4/3s=4/3 therefore s=1